OpenStudy (anonymous):

Find the surface area for a cylinder with a radius of 4 and a height of 6. (use 3.14 for pi) Answer 251.20 150.72 200.96 156.78

OpenStudy (anonymous):

\[2\pi *r*h=150.72\]

OpenStudy (anonymous):

Using the formula for the surface are of a cylinder: 2(pi)*r^2 + 2(pi)*r*h , where r=4 and h=6: is 251.2